# PrismAudio: Decomposed Chain-of-Thoughts and Multi-dimensional Rewards for Video-to-Audio Generation

### v3

@@ F.1 Objective Evaluation
Semantic Consistency: We utilize the CLAP (Contrastive Language-Audio Pre-training) score (elizalde2024natural) to measure semantic alignment between generated audio and our constructed Chain-of-Thought descriptions in a shared audio-text embedding space. The CLAP score provides a robust measure of how well the generated audio semantically matches the detailed reasoning and content descriptions provided by our CoT framework.
Audio-Visual Synchrony: To evaluate temporal synchronization between generated audio and corresponding video, we adopt the DeSync score predicted by the Synchformer model (iashin2024synchformer). For each sample, we truncate the video to match the duration of the generated audio and compute the DeSync score using Synchformer’s 4.8-second context window. Specifically, we extract both the first and last 4.8-second segments from each video-audio pair, calculate DeSync scores for each segment, and report the average as the final temporal alignment metric. Lower DeSync scores indicate better synchronization.
Aesthetic Quality: We employ four complementary metrics from Audiobox-Aesthetics (tjandra2025meta) to comprehensively assess aesthetic quality:
@@
  • GCC MSE: Utilizes the traditional Generalized Cross-Correlation with Phase Transform (GCC-PHAT) (knapp2003generalized) to estimate TDOA between left and right channels. We compute the mean squared error between ground truth and generated audio TDOA distributions.
  • CRW MSE: Employs the deep learning-based StereoCRW network (chen2022sound) for TDOA estimation. Similar to GCC MSE, we calculate the mean squared error between reference and generated TDOA distributions.
Feature Distribution Alignment: We further assess the similarity to real audio distributions using two reference metrics. The Fréchet Distance (FD) on VGGish embeddings (kilgour2018fr) measures statistical realism, while the Kullback-Leibler (KL) Divergence (copet2024simple) on PaSST classifier (koutini2021efficient) outputs evaluates content plausibility. Crucially, both metrics act as imperfect proxies, as they rely on fixed, pre-trained models and do not capture the fine-grained, conditional alignment (e.g., temporal, spatial) that is central to our evaluation. Thus, they serve as valuable supplementary indicators to gauge overall audio quality, rather than as primary measures of performance.
@@ Appendix G Limitation and Future Work
While PrismAudio successfully establishes a new paradigm for video-to-audio generation that effectively resolves objective entanglement, its current implementation highlights several boundaries and exciting opportunities for future research. (1) Our current decoupled paradigm, while effective, relies heavily on the capabilities of the upstream MLLMs planner. This creates a “cascading error" problem: any misinterpretations by the planner are irreversibly passed down to the audio generator, which can only optimize for a potentially flawed plan. A significant leap forward would be to explore unified, end-to-end architectures that jointly learn to perceive, reason, and generate within a single model. Such a model could mitigate the error propagation issue by allowing for a more deeply integrated reasoning and synthesis process, potentially leading to more coherent and grounded results. (2) A second promising direction involves advancing the multi-dimensional reinforcement learning stage. Our current framework aggregates the four reward signals using a static weighting policy, applying the same balance of priorities to all videos. However, the perceptual importance of each dimension can vary significantly with video content; for example, a fast-paced action sequence demands prioritizing precise temporal synchrony, whereas a tranquil landscape benefits more from high aesthetic fidelity. A major advancement would be to develop a content-aware RL policy. Such a system could learn to dynamically adjust the weights of the different reward functions based on the input video’s content.
@@ Appendix H Ethical Considerations
  • Bias and Representation Issues. We acknowledge that our training data and reward functions may inadvertently encode cultural biases regarding what constitutes “good sound" or appropriate audio-visual relationships. The Aesthetic CoT module and corresponding reward signals could reflect specific perceptual preferences that may not generalize across diverse cultural contexts, potentially marginalizing non-Western audio traditions or alternative aesthetic standards. However, we have implemented several significant practical mitigations in our work to address these concerns:
    (1) Technical Quality Focus: A crucial aspect of our framework is how we define and operationalize “aesthetics." This is not a vague, culturally-loaded preference. As stated in our paper, our Aesthetic CoT focuses on tangible audio quality aspects like “naturalness and fidelity." This is reinforced by our choice of reward model. As detailed in Appendix F.1, the Meta Audiobox Aesthetics model provides multi-faceted Aesthetics scores. While it includes a subjective “Content Enjoyment (CE)" score, our optimization also heavily relies on the “Production Quality (PQ)" metric. PQ explicitly measures technical aspects like clarity, fidelity, dynamics, and frequency response, rather than subjective tastes. By optimizing for high fidelity and technical clarity, we primarily push the model towards realism and professional production standards, which are far more universal and less culturally specific than artistic or stylistic preferences. This anchors our “aesthetic" goal to a more objective standard, reducing the risk of encoding culturally-biased preferences.
@@
    While we do not claim to have eliminated all bias, we believe our specific methodological choices—from the scope of our task to the definition and measurement of “aesthetics"—serve as significant, practical mitigation steps. These choices reflect our commitment to responsible AI development and demonstrate that bias mitigation can be integrated into the core design of the system, rather than being treated as an afterthought. Future work should continue to prioritize inclusive dataset construction, culturally aware reward design, and ongoing evaluation of potential biases in generated content.
@@ Appendix I Potential Negative Societal Impacts
We disclose that LLMs were used during the preparation of this manuscript. We utilized LLMs exclusively as a writing aid to enhance the clarity and readability of the text.

### v4

@@ F.1 Objective Evaluation
Semantic Consistency: We utilize the CLAP (Contrastive Language-Audio Pre-training) score (Elizalde et al., 2024) to measure semantic alignment between generated audio and our constructed Chain-of-Thought descriptions in a shared audio-text embedding space. The CLAP score provides a robust measure of how well the generated audio semantically matches the detailed reasoning and content descriptions provided by our CoT framework.
Audio-Visual Synchrony: To evaluate temporal synchronization between generated audio and corresponding video, we adopt the DeSync score predicted by the Synchformer model (Iashin et al., 2024). For each sample, we truncate the video to match the duration of the generated audio and compute the DeSync score using Synchformer’s 4.8-second context window. Specifically, we extract both the first and last 4.8-second segments from each video-audio pair, calculate DeSync scores for each segment, and report the average as the final temporal alignment metric. Lower DeSync scores indicate better synchronization.
Aesthetic Quality: We employ four complementary metrics from Audiobox-Aesthetics (Tjandra et al., 2025) to comprehensively assess aesthetic quality:
@@
  • GCC MSE: Utilizes the traditional Generalized Cross-Correlation with Phase Transform (GCC-PHAT) (Knapp and Carter, 2003) to estimate TDOA between left and right channels. We compute the mean squared error between ground truth and generated audio TDOA distributions.
  • CRW MSE: Employs the deep learning-based StereoCRW network (Chen et al., 2022) for TDOA estimation. Similar to GCC MSE, we calculate the mean squared error between reference and generated TDOA distributions.
Feature Distribution Alignment: We further assess the similarity to real audio distributions using two reference metrics. The Fréchet Distance (FD) on VGGish embeddings (Kilgour et al., 2018) measures statistical realism, while the Kullback-Leibler (KL) Divergence (Copet et al., 2024) on PaSST classifier (Koutini et al., 2021) outputs evaluates content plausibility. Crucially, both metrics act as imperfect proxies, as they rely on fixed, pre-trained models and do not capture the fine-grained, conditional alignment (e.g., temporal, spatial) that is central to our evaluation. Thus, they serve as valuable supplementary indicators to gauge overall audio quality, rather than as primary measures of performance.
@@ Appendix G Limitation and Future Work
While PrismAudio successfully establishes a new paradigm for video-to-audio generation that effectively resolves objective entanglement, its current implementation highlights several boundaries and exciting opportunities for future research. (1) Our current decoupled paradigm, while effective, relies heavily on the capabilities of the upstream MLLMs planner. This creates a “cascading error” problem: any misinterpretations by the planner are irreversibly passed down to the audio generator, which can only optimize for a potentially flawed plan. A significant leap forward would be to explore unified, end-to-end architectures that jointly learn to perceive, reason, and generate within a single model. Such a model could mitigate the error propagation issue by allowing for a more deeply integrated reasoning and synthesis process, potentially leading to more coherent and grounded results. (2) A second promising direction involves advancing the multi-dimensional reinforcement learning stage. Our current framework aggregates the four reward signals using a static weighting policy, applying the same balance of priorities to all videos. However, the perceptual importance of each dimension can vary significantly with video content; for example, a fast-paced action sequence demands prioritizing precise temporal synchrony, whereas a tranquil landscape benefits more from high aesthetic fidelity. A major advancement would be to develop a content-aware RL policy. Such a system could learn to dynamically adjust the weights of the different reward functions based on the input video’s content.
@@ Appendix H Ethical Considerations
  • Bias and Representation Issues. We acknowledge that our training data and reward functions may inadvertently encode cultural biases regarding what constitutes “good sound” or appropriate audio-visual relationships. The Aesthetic CoT module and corresponding reward signals could reflect specific perceptual preferences that may not generalize across diverse cultural contexts, potentially marginalizing non-Western audio traditions or alternative aesthetic standards. However, we have implemented several significant practical mitigations in our work to address these concerns:
    (1) Technical Quality Focus: A crucial aspect of our framework is how we define and operationalize “aesthetics.” This is not a vague, culturally-loaded preference. As stated in our paper, our Aesthetic CoT focuses on tangible audio quality aspects like “naturalness and fidelity.” This is reinforced by our choice of reward model. As detailed in Appendix F.1, the Meta Audiobox Aesthetics model provides multi-faceted Aesthetics scores. While it includes a subjective “Content Enjoyment (CE)” score, our optimization also heavily relies on the “Production Quality (PQ)” metric. PQ explicitly measures technical aspects like clarity, fidelity, dynamics, and frequency response, rather than subjective tastes. By optimizing for high fidelity and technical clarity, we primarily push the model towards realism and professional production standards, which are far more universal and less culturally specific than artistic or stylistic preferences. This anchors our “aesthetic” goal to a more objective standard, reducing the risk of encoding culturally-biased preferences.
@@
    While we do not claim to have eliminated all bias, we believe our specific methodological choices—from the scope of our task to the definition and measurement of “aesthetics”—serve as significant, practical mitigation steps. These choices reflect our commitment to responsible AI development and demonstrate that bias mitigation can be integrated into the core design of the system, rather than being treated as an afterthought. Future work should continue to prioritize inclusive dataset construction, culturally aware reward design, and ongoing evaluation of potential biases in generated content.
@@ Appendix I Potential Negative Societal Impacts
In accordance with ICLR’s guidelines, we disclose that LLMs were used during the preparation of this manuscript. We utilized LLMs exclusively as a writing aid to enhance the clarity and readability of the text.